# Adaptive Incentive Design with Learning Agents

### v2

@@ 4.2 Non-atomic Traffic Routing on General Networks
Condition (44) ensures that the function $\tilde{\phi}$ is differentiable in a neighborhood of $\tilde{p}^{\dagger}$ (cf. [Yang and Huang, 2005, Chapter 4]). From Lemma B.1, we know that if $z^{\top}\Delta D\phi(\tilde{p}^{\dagger})z\leqslant 0$ for all $z$ then there exists $r>0$ such that (46) holds for all $\tilde{p}\in\mathcal{B}_{r}(\tilde{p}^{\dagger})$. Indeed, by design of $\Delta$ it holds that $\Delta D\phi(\tilde{p}^{\dagger})=D\tilde{w}(\tilde{p}^{\dagger})$. Finally, using Lemma B.1 and Lemma B.4, we note that $z^{\top}D\tilde{w}(\tilde{p}^{\dagger})z\leqslant 0$. This concludes the proof. ∎
5 Concluding Remarks
We propose an adaptive incentive mechanism that $(i)$ updates based on the feedback of each agent’s externality, $(ii)$ is agnostic to the learning rule employed by agents, and $(iii)$ updates at a slower timescale compared to the agents’ learning dynamics, leading to a two-timescale coupled dynamical system. We provide sufficient conditions that ensure the convergence of this dynamical system to the optimal incentive mechanism, guaranteeing that the Nash equilibrium of the corresponding game among agents achieves social optimality. We demonstrate the effectiveness of our proposed incentive mechanism in two practically relevant games: atomic networked quadratic aggregative games and non-atomic network routing games, by verifying the proposed convergence conditions.
An important direction for future research is to verify the convergence conditions in a broader class of games. Notably, leveraging the externality-based dynamic incentive design update proposed in an earlier version of this paper (cf. [Maheshwari et al., 2021]), [Li et al., 2023a] studied the convergence of such updates in power systems.
Acknowledgements
Appendix A Counter-example.
In this section, we present an instance of a non-atomic game where the standard gradient-based incentive design approach would fail. Specifically, we will show that the social cost, as a function of the incentive (evaluated at Nash equilibrium), is non-convex and has a zero gradient, which would result in the gradient-based approach not converging.
Figure 1: Two-link routing game.
Thus, for any toll vector $\tilde{p}$ the Nash equilibrium is ${\tilde{x}^{\ast}}(\tilde{p})=\arg\min_{\tilde{x}\in\tilde{X}}T(\tilde{x},\tilde{p})$ is such that ${\tilde{x}^{\ast}_{1}}(\tilde{p})=\mathcal{P}_{[0,1]}\left(\frac{\tilde{p}_{2}-\tilde{p}_{1}+1}{2}\right),{\tilde{x}^{\ast}_{2}}(\tilde{p})=\mathcal{P}_{[0,1]}\left(\frac{\tilde{p}_{1}-\tilde{p}_{2}+1}{2}\right),$ where for any scalar $x\in\mathbb{R}$, $\mathcal{P}_{[0,1]}(x)$ denotes its projection on the line segment $[0,1]$.
The social cost function (cf. (36)) is given by $\tilde{\Phi}(\tilde{x})=\tilde{x}_{1}\tilde{\ell}_{1}(\tilde{x})+\tilde{x}_{2}\tilde{\ell}_{2}(\tilde{x})=\tilde{x}_{1}^{2}+\tilde{x}_{2}^{2}$. Thus,
$\tilde{\Phi}({\tilde{x}^{\ast}}(\tilde{p}))$ =$=\begin{cases}\frac{(\tilde{p}_{1}-\tilde{p}_{2})^{2}+1}{2}&\text{ if }|\tilde{p}_{1}-\tilde{p}_{2}|\leqslant 1,\\ 1&\text{ otherwise }\end{cases}$
However, this function is not convex, for example, on the set $|\tilde{p}_{1}-\tilde{p}_{2}|<1.5$. Therefore, gradient based updates initialized in the region $\{(\tilde{p}_{1},\tilde{p}_{2})\in\mathbb{R}^{2}:1.1\leqslant|\tilde{p}_{1}-\tilde{p}_{2}|\leqslant 1.5\}$ will not converge to the global optimizer as the gradient is zero in this region. Thus, the approach from [Li et al., 2023b, Liu et al., 2021] does not apply here.
Appendix B Auxiliary Results
Lemma B.1 ([Facchinei and Pang, 2007]).
For any fixed $p^{\prime}$ and continuously differentiable function ϕ:~~:italic-$\phi:\mathbb{R}^{\tilde{\mathcal{E}}}\rightarrow\mathbb{R}^{\tilde{\mathcal{E}}}$ the condition
for some $r>0$, holds if and only if $z^{\top}D\phi(p^{\prime})z\leqslant 0$ for all $z\in\mathbb{R}^{|\tilde{\mathcal{E}}|}$.
Lemma B.2.
The externality $\tilde{e}_{i}^{j}$ in (37) can be alternatively written, in terms of edge flows, as $\tilde{e}_{i}^{j}(\tilde{x})=\sum_{a\in j}\tilde{w}_{a}\nabla l_{a}(\tilde{w}_{a}).$
@@ Appendix B Auxiliary Results
Expanding the externality expression from (37) we observe that
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{i^{\prime}\in\tilde{\mathcal{I}}}\sum_{j^{\prime}\in\tilde{R}_{i}}\tilde{x}_{i^{\prime}}^{j^{\prime}}\sum_{a\in\tilde{\mathcal{E}}}\mathbbm{1}(a\in j^{\prime})\frac{\partial l_{a}(\tilde{w}_{a})}{\partial\tilde{x}_{i}^{j}}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{i^{\prime}\in\tilde{\mathcal{I}}}\sum_{j^{\prime}\in\tilde{R}_{i}}\tilde{x}_{i^{\prime}}^{j^{\prime}}\sum_{a\in\tilde{\mathcal{E}}}\mathbbm{1}(a\in j^{\prime})\nabla l_{a}(\tilde{w}_{a})\frac{\partial\tilde{w}_{a}}{\partial\tilde{x}_{i}^{j}}$
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{i^{\prime}\in\tilde{\mathcal{I}}}\sum_{j^{\prime}\in\tilde{R}_{i}}\tilde{x}_{i^{\prime}}^{j^{\prime}}\sum_{a\in\tilde{\mathcal{E}}}\mathbbm{1}(a\in j^{\prime})\nabla l_{a}(\tilde{w}_{a})\mathbbm{1}(a\in j)$
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{a\in\tilde{\mathcal{E}}}\nabla l_{a}(\tilde{w}_{a})\mathbbm{1}(a\in j)\sum_{i^{\prime}\in\tilde{\mathcal{I}}}\sum_{j^{\prime}\in\tilde{R}_{i}}\tilde{x}_{i^{\prime}}^{j^{\prime}}\mathbbm{1}(a\in j^{\prime})$
=$\stackrel{{\scriptstyle(e)}}{{=}}\sum_{a\in j}\nabla l_{a}(\tilde{w}_{a})\tilde{w}_{a}$
where $(a)$ follows by expanding out the expression of route costs in terms of edge costs, $(b)$ follows by the chain rule, $(c)$ follows due to the definition of edge flows, $(d)$ follows by change of order of summations and finally $(e)$ follows by the definition of edge flows. This completes the proof. ∎
Lemma B.3.
@@
First, we show that
$\Phi(\tilde{x})=\sum_{a\in\tilde{\mathcal{E}}}\tilde{w}_{a}l_{a}(\tilde{w}_{a}).$ (47)
Indeed,
@@
$=\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\tilde{x}_{i}^{j}\sum_{a\in\tilde{\mathcal{E}}}\mathbbm{1}(a\in j)l_{a}(\tilde{w}_{a})$
@@
$=\sum_{a\in\tilde{\mathcal{E}}}\frac{\partial}{\partial\tilde{w}_{a}}\left(\tilde{w}_{a}l_{a}(\tilde{w}_{a})\right)\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}\mathbbm{1}(a\in j)(\tilde{x}_{i}^{j}-\tilde{x}_{i}^{\dagger j})$
@@
Lemma B.4.
For any $\tilde{p},\tilde{p}^{\prime}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}$ the following holds
$\sum_{a\in\tilde{\mathcal{E}}}(\tilde{p}_{a}-\tilde{p}^{\prime}_{a})(\tilde{w}^{\ast}_{a}(\tilde{p}_{a})-\tilde{w}^{\ast}_{a}(\tilde{p}^{\prime}_{a}))\leqslant 0.$ (48)
@@
where $\tilde{P}$ and $\tilde{P}^{\prime}$ are the route tolls associated with edge tolls $\tilde{p}$ and $\tilde{p}^{\prime}$ respectively through (35).
Let the feasible set of the optimization problem (39) be denoted by $\mathcal{F}$. Using the first-order conditions of optimality for the strictly convex optimization problem (39) we obtain the following variational inequality:
@@
We are now ready to show the condition (49). Note that
@@
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{a\in\tilde{\mathcal{E}}}(l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\prime}))-l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p})))(\tilde{w}^{\ast}_{a}(\tilde{p})-\tilde{w}^{\ast}_{a}(\tilde{p}^{\prime}))$
$\stackrel{{\scriptstyle(e)}}{{\leqslant}}0$ (52)
where we obtain $(a)$ by adding (50), evaluated at $\tilde{y}=\tilde{x}^{\ast}(\tilde{P}^{\prime})$, and (51), evaluated at $\tilde{y}^{\prime}=\tilde{x}^{\ast}(\tilde{P})$, $(b)$ holds by definition of route loss function, $(c)$ holds by interchange of summation, $(d)$ holds by definition of edge flows and $(e)$ holds due to monotonicity of edge latency functions. This proves (49).
Now we are ready to prove (48). Note that
@@
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}(\tilde{x}_{i}^{\ast j}(\tilde{P})-\tilde{x}_{i}^{\ast j}(\tilde{P}^{\prime}))(\tilde{P}_{i}^{j}-\tilde{P}_{i}^{{}^{\prime}j})$
$\stackrel{{\scriptstyle(d)}}{{\leqslant}}0,$
where $(a)$ holds due to the definition of edge flows, $(b)$ holds due to interchange of summation, $(c)$ holds due to definition of route tolls, $(d)$ holds due to (49). This concludes the proof. ∎

### v3

@@ B.3 Additional Results
To show this, we note that due to condition (36), the function $\phi$ is differentiable in a neighborhood of $\tilde{p}^{\dagger}$ (cf. [Yang and Huang, 2005, Chapter 4]). Consequently, using Lemma D.5 in Appendix D, it is sufficient to show that
$\sum_{a,a^{\prime}\in\tilde{\mathcal{E}}}z_{a}\Delta_{a,a}\frac{\partial\phi_{a}(\tilde{p}^{\dagger})}{\partial\tilde{p}_{a^{\prime}}}z_{a^{\prime}}\leqslant 0,\quad\forall\ z\in\mathbb{R}^{|\tilde{\mathcal{E}}|}.$ (57)
Indeed, by the design of $\Delta$, it holds that
$\Delta_{a,a}\frac{\partial\phi_{a}(\tilde{p}^{\dagger})}{\partial\tilde{p}_{a^{\prime}}}=\frac{\partial\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})}{\partial\tilde{p}_{a^{\prime}}},\quad\forall\ a,a^{\prime}\in\tilde{\mathcal{E}}.$ (58)
Furthermore, Lemma D.4 and Lemma D.5 in Appendix D guarantee that
$\sum_{a,a^{\prime}\in\tilde{\mathcal{E}}}z_{a}\frac{\partial\tilde{w}^{\ast}_{a}(\tilde{p}^{\dagger})}{\partial\tilde{p}_{a^{\prime}}}z_{a^{\prime}}\leqslant 0,\quad\forall\ z\in\mathbb{R}^{|\tilde{\mathcal{E}}|}.$ (59)
The proof concludes by noting that (58) and (59) imply (57).
Appendix D Auxiliary Results
Lemma D.1.
Requirement (R1) of Proposition 3.3 is satisfied if either one of the following conditions holds:
  • (C1)
    $\frac{\partial e_{i}(x^{\ast}(p))}{\partial p_{j}}>0$ for all $p\in\mathbb{R}^{n}$ and all $i\neq j$, and at least one of the following conditions holds:
    • (i)
      $e_{i}(x^{\ast}(0))\geqslant 0$ for every $i\in\mathcal{I}$, $p^{\dagger}\in\mathbb{R}_{+}^{|\mathcal{I}|},$ and for any $p\in\mathbb{R}_{+}^{|\mathcal{I}|}$, there exists $p^{\prime}\in\mathbb{R}_{+}^{|\mathcal{I}|}$ such that $p^{\prime}_{i}>p_{i}$ and $e_{i}(x^{\ast}(p^{\prime}))-p_{i}^{\prime}\leqslant 0~{}\text{for every }i\in\mathcal{I}.$ Moreover, $x_{0}\in X,p_{0}\in\mathbb{R}^{|\mathcal{I}|}_{+}$.
    • (ii)
      $e_{i}(x^{\ast}(0))\leqslant 0$ for every $i\in\mathcal{I}$, $p^{\dagger}\in\mathbb{R}_{-}^{|\mathcal{I}|},$ and for any $p\in\mathbb{R}_{-}^{|\mathcal{I}|}$, there exists $p^{\prime}\in\mathbb{R}_{-}^{|\mathcal{I}|}$ such that $p^{\prime}_{i}<p_{i}$ and $e_{i}(x^{\ast}(p^{\prime}))-p_{i}^{\prime}\geqslant 0~{}\text{for every }i\in\mathcal{I}.$ Moreover, $x_{0}\in X,p_{0}\in\mathbb{R}^{|\mathcal{I}|}_{-}$.
  • (C2)
    There exists a set $\textsf{dom}(V)\subset\mathbb{R}^{|\mathcal{I}|}$ and a continuously differentiable function $V:\textsf{dom}(V)\rightarrow\mathbb{R}_{+}$ such that $V(p^{\dagger})=0$ and $V(p)>0$ for all $p\neq p^{\dagger}$. Moreover, for every $p\neq p^{\dagger},$ $\nabla V(p)^{\top}\left(e(x^{\ast}(p))-p\right)<0.$
@@ Appendix D Auxiliary Results
Conditions (C1) and (C2) above are based on results from non-linear dynamical systems which ensure convergence of (19). In particular, (C1)-(i) (resp. (C1)-(ii)) builds on cooperative dynamical systems theory [Hirsch, 1985], which ensures that $\mathbb{R}_{+}^{|\mathcal{I}|}$ (resp. $\mathbb{R}_{-}^{|\mathcal{I}|}$) is positively invariant for (19) and $p^{\dagger}\in\mathbb{R}_{+}^{|\mathcal{I}|}$ (resp. $p^{\dagger}\in\mathbb{R}_{-}^{|\mathcal{I}|}$) is asymptotically stable. On the other hand, condition (C2) ensures the existence of a Lyapunov function that is strictly positive everywhere except at $p^{\dagger}$ and decreases along any trajectory of (19) (cf. [Sastry, 2013]). ∎
Lemma D.2.
For every $i\in\tilde{\mathcal{I}},j\in\tilde{R}_{i},$ $\tilde{e}_{i}^{j}(\tilde{x})=\sum_{a\in j}\tilde{w}_{a}\nabla l_{a}(\tilde{w}_{a}).$
Proof.
Using (34), we note that
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{i^{\prime}\in\tilde{\mathcal{I}}}\sum_{j^{\prime}\in\tilde{R}_{i}}\tilde{x}_{i^{\prime}}^{j^{\prime}}\sum_{a\in\tilde{\mathcal{E}}}\mathbbm{1}(a\in j^{\prime})\nabla l_{a}(\tilde{w}_{a})\frac{\partial\tilde{w}_{a}}{\partial\tilde{x}_{i}^{j}}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{i^{\prime}\in\tilde{\mathcal{I}}}\sum_{j^{\prime}\in\tilde{R}_{i}}\tilde{x}_{i^{\prime}}^{j^{\prime}}\sum_{a\in\tilde{\mathcal{E}}}\mathbbm{1}(a\in j^{\prime})\nabla l_{a}(\tilde{w}_{a})\mathbbm{1}(a\in j)$
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{a\in j}\nabla l_{a}(\tilde{w}_{a})\tilde{w}_{a},$
where $(a)$ follows by expanding out the expression of route costs in terms of edge costs and using the chain rule, $(b)$ follows by the definition of edge flows, and $(c)$ follows by changing the order of summations and using the definition of edge flows. This completes the proof. ∎
Lemma D.3.
@@
First, we show that $\Phi(\tilde{x})=\sum_{a\in\tilde{\mathcal{E}}}\tilde{w}_{a}l_{a}(\tilde{w}_{a}).$
@@
Lemma D.4.
Following inequality holds:
$\sum_{a\in\tilde{\mathcal{E}}}(\tilde{p}_{a}-\tilde{p}^{\prime}_{a})\left(\tilde{w}^{\ast}_{a}(\tilde{p}_{a})-\tilde{w}^{\ast}_{a}(\tilde{p}^{\prime}_{a})\right)\leqslant 0,\quad\forall\ \tilde{p},\tilde{p}^{\prime}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}.$ (60)
@@
where $\tilde{P}$ and $\tilde{P}^{\prime}$ are the route tolls associated with edge tolls $\tilde{p}$ and $\tilde{p}^{\prime}$, respectively, through (32). Let the feasible set of route flows in the optimization problem (51) be denoted by $\mathcal{F}$. Using the first-order conditions of optimality for the strictly convex optimization problem (51), we obtain:
@@
Next, we prove (61). Note that
@@
=$\stackrel{{\scriptstyle(d)}}{{=}}\sum_{a\in\tilde{\mathcal{E}}}(l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p}^{\prime}))-l_{a}(\tilde{w}^{\ast}_{a}(\tilde{p})))(\tilde{w}^{\ast}_{a}(\tilde{p})-\tilde{w}^{\ast}_{a}(\tilde{p}^{\prime}))\stackrel{{\scriptstyle(e)}}{{\leqslant}}0,$
where we obtain $(a)$ by adding (62), evaluated at $\tilde{y}=\tilde{x}^{\ast}(\tilde{P}^{\prime})$, and (63), evaluated at $\tilde{y}^{\prime}=\tilde{x}^{\ast}(\tilde{P})$, $(b)$ holds by the definition of the route loss function, $(c)$ holds by interchange of summation, $(d)$ holds by the definition of edge flows, and $(e)$ holds due to the monotonicity of edge latency functions. This proves (61).
Finally, we prove (60). Note that
@@
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{i\in\tilde{\mathcal{I}}}\sum_{j\in\tilde{R}_{i}}(\tilde{x}_{i}^{\ast j}(\tilde{P})-\tilde{x}_{i}^{\ast j}(\tilde{P}^{\prime}))(\tilde{P}_{i}^{j}-\tilde{P}_{i}^{{}^{\prime}j})\stackrel{{\scriptstyle(d)}}{{\leqslant}}0,$
where $(a)$ holds due to the definition of edge flows, $(b)$ holds due to interchange of summation, $(c)$ holds due to the definition of route tolls, and $(d)$ holds due to (61). This concludes the proof. ∎
Lemma D.5 ([Facchinei and Pang, 2007]).
For any fixed $p^{\prime}$ and continuously differentiable function ϕ:~~:italic-$\phi:\mathbb{R}^{\tilde{\mathcal{E}}}\rightarrow\mathbb{R}^{\tilde{\mathcal{E}}}$, the condition
for some $r>0$, holds if and only if
$\sum_{i,j\in|\tilde{\mathcal{E}}|}z_{i}z_{j}\frac{\partial\phi_{i}(p^{\prime})}{\partial p_{j}}\leqslant 0,\quad\forall\ z\in\mathbb{R}^{|\tilde{\mathcal{E}}|}.$